Show that there does not exist a non-constant continuous function f : R2 → R such that f(x, y) = 5 for all (x, y) ∈ R2 with x^2+y^2<1


1
Expert's answer
2021-01-04T20:48:22-0500

Let the function

"f:\R^2\rightarrow \R\ni\\
f(x,y) = \begin{cases}
 5 &\text{if } x^2+y^2<1 \\
 x^2+y^2+4 &\text{if } x^2+y^2\geq1 
\end{cases}"


We can see the continuity of "f(x,y)" by two processes.

(i)

for "x^2+y^2<1" , "f(x,y)" is a constant function.


for "x^2+y^2<1 , f(x,y)= x^2+y^2+4" which is sum of three continuous functions,"\;x^2,y^2\;and \;4." So it is continuous.


for "x^2+y^2=1" , "\lim_{\{(x,y)|x^2+y^2\rightarrow 1\}}f(x,y)=5" from all sides.


Therefore we can conclude that, "f(x,y)\; is \; continuous \;\forall(x,y)\in \R^2" .


(ii) [By "\epsilon-\delta" definition]

Let a point be "(h,k)"

Given "\epsilon>0,"


If "\; h^2+k^2<1 ,"

Take "\;\delta = min\{\sqrt{ h^2+k^2},1-\sqrt{ h^2+k^2}\}"

Let "(x,y)\in\R^2\ni d((h,k),(x,y))<\delta"

"\Rightarrow f(x,y)=5\\
\Rightarrow |f(x,y)-f(h,k)|=|5-5|=0<\epsilon"


If "h^2+k^2\geq1,"

Take "\delta =\sqrt{ h^2+k^2}-\sqrt{ h^2+k^2-\epsilon}"

Let "(x,y)\in\R^2\ni d((h,k),(x,y))<\delta"

"\Rightarrow (x,y)\in \{(p,q)|(x-h)^2+(y-k)^2<\delta^2\}\\
\Rightarrow f(x,y)\in \left[f(p_1,q_1),f(p_2,q_2)\right]\\
where,\; p_1^2+q_1^2=(\sqrt{h^2+k^2}-\delta)^2=(\sqrt{h^2+k^2-\epsilon})^2=h^2+k^2-\epsilon\\
and, p_2^2+q_2^2=(\sqrt{h^2+k^2}+\delta)^2\leq h^2+k^2+\epsilon\\
\Rightarrow f(x,y)\in [h^2+k^2+4-\epsilon,h^2+k^2+4+\epsilon]\\
\Rightarrow f(x,y)\in [f(h,k)-\epsilon,f(h,k)+\epsilon]\\
\Rightarrow |f(x,y)-f(h,k)|<\epsilon"


So, "d((h,k),(x,y))<\delta \Rightarrow d(f(x,y),f(h,k))<\epsilon"


Hence "\pmb{f(x,y)}" is continuous "\pmb{\forall(x,y)\in \R^2}" .